A university graduate standing on the ground fires a rocket into the atmosphere at a velocity of 4000 m/s at an angle of 20 degrees to the vertical axis. Determine (g = 9.8 m/s2):


(a) The time of flight

(b) The maximum level


1
Expert's answer
2022-03-27T18:21:20-0400

"t=\frac{2v\cos\alpha}g=12.8~min,"

"h=\frac{(v\cos\alpha)^2}{2g}=721~km."